# Equality Saturation for Optimizing High-Level Julia IR

### v1

@@ 4.3. Multi-Line Broadcast Fusion
Julia code that uses broadcasting syntax: using the dot operator, a elementwise ReLu operation and an elementwise addition are applied to vectors A and B. In other words, the relu function is broadcasted over matrix A and vector B is added to the result using broadcasting addition.
@@
Julia’s broadcasting mechanism allows users to apply functions element-by-element on one or more arguments containing multiple elements. Syntactically, this is done by adding a . between the function and its argument list. The top of Figure 7 shows two broadcast operations. relu.(A) applies the scalar relu function to each element of the matrix A. Next, the .+ call computes the addition between the result and a vector B. Although the two arguments to .+ have a different shape, the operation can take place because the shapes are compatible.
Broadcasting in Julia is implemented by transforming expressions containing this special dot syntax into calls to the standard library functions broadcasted and materialize. Figure 8 shows the result of this transformation on the broadcasting expression in Figure 7. A call to broadcasted builds a lazy representation of the broadcast result. Different lazy broadcast objects can be nested; it is only when materialize is called that the final result object is allocated and filled with computed values. As such, different operations within multiple broadcasted calls are fused when materialize is called. This leads to less memory being used because temporary arrays are not materialized and potentially better run-time performance because there are fewer function calls. This operator fusion is especially beneficial for code that is executed on the GPU, as kernel launches can take a significant amount of time, and because kernel fusion can significantly reduce the required number of memory accesses (Besard et al., 2019).
A limitation of Julia’s broadcasting lowering is that fusion cannot occur across broadcast expressions on different lines. This is because replacing the dot syntax with calls to broadcasted and materialize is a syntactical transformation that is performed during AST lowering, which operates statement by statement. A potential optimization is hence to eliminate superfluous calls to materialize. Rewrite rules make this easy. Take for example the rule
This rule will remove the intermediary materialization for any call of the form f.(x) where x is a variable that was defined on a different line with a broadcast expression itself. In the example code in Figure 9, fusion will now be applied despite the two dot operators occurring in two different statements on two separate lines. This example illustrates how the system we propose not only supports domain-specific optimization, but at the same time enables optimization of common code patterns involving only core Julia primitives.
An important detail is that here, once more, CFG skeleton relaxation is required, since materialize ends up calling a foreign C function function which the compiler cannot deem effect-free, poisoning the remainder of the effect analysis. In practice, this means the rule needs to be prefixed with unsafe.
@@
For other patterns, e.g., an expression where the outer broadcasted function is a binary function, separate rules need to be written. Table 1 gives an overview of how many times different multi-line broadcasting patterns were found in our analysis. The three last patterns are different configurations of broadcasting a binary function. Note that the matches for these three patterns are not independent: a match against the last pattern, which will result in two invocations of materialize being optimized out, implies the two other patterns can be matched and applied as well, each optimizing out only one materialize invocation. When the final code is extracted in such cases, the expression from the e-class containing the least materialize calls will be chosen.
@@ 4.4. Extending Broadcasting with Domain-Specific Batched Operations
In high-performance contexts, programmers often sacrifice code simplicity and readability in pursuit of achieving the highest possible performance. For example, instead of writing deep learning layers and non-linear activation functions using core language constructs, programmers often resort to using frameworks or libraries that contain optimized implementations of matrix multiplications, activation functions, attention computation, and others (Elshawi et al., 2021; Parvat et al., 2017). This works well for most common use cases but falls apart when a programmer requires custom behavior that is not yet supported by the framework. In those cases, the programmer is required to rewrite parts of the program to stop using the framework and instead depend on a different framework or the core language.
Instead of having to opt out of framework functionality and rewriting code, it would be much better if one could write simple, readable code and have the frameworks themselves automatically discover opportunities to use their better performing implementations. Our rewrite system gives framework authors the option for implementing such discovery and optimization almost for free.
@@
As an example, take the code in Figure 9 that shows an implementation of the forward pass of a single fully connected feedforward layer. There already exist several Julia packages that allow us to compute such operation more efficiently (Faingnaert et al., 2022; Innes et al., 2018; Innes, 2018). For example, the programmer could rewrite the code as shown in Figure10 to use the optimized fused_dense_bias_activation function from the Lux.jl deep learning framework (Pal, 2023). A disadvantage of such manual rewriting is that the code becomes less readable for someone not familiar with the Lux.jl APIs. Moreover, the use of that specific framework’s optimized implementation, which is in essence a decision about a non-functional aspect of the code, is then hard-coded in the code that specifies the functionality. This requires the developers and the maintainers of this code to be knowledgable in the code’s application domain as well as in the Lux.jl framework. Switching to other frameworks in the future will require rewriting code. Clearly, hard coding the dependence on Lux.jl and its APIs in the code has a number of downsides.
@@
With our system, instead of rewriting the original code of Figure 9, it suffices to add a rewrite rule that maps the original code pattern onto the optimized fused_dense_bias_activation function:999Type annotations have been left out for the sake of conciseness of this rule, but they are needed to make sure to only match when W and x are matrices, and b is a vector.
| materialize(broadcasted(~, broadcasted(+, ~W * ~x, ~b))) | ||
@@
Combined with the multi-line broadcast rewriting rule explained above, which rewrites the original code into a single line without the intermediary variable logits, this rule maps the single-line broadcast expression onto a call to the library function LuxLib.fused_dense_bias_activation, which will yield compiled code equivalent to what would be obtained with the code in Figure 10.
@@
This example illustrates how our approach allows for a better separation of concerns. Another advantage is that alternative rewrite rules can be provided for different Julia packages that provide different optimized implementations for similar functionality. Different rewrite rules also don’t obfuscate each other. All rewrite rules can be represented in the e-graph. Only during extraction is the final optimized code materialized, optionally based on user-defined extraction costs for different functions. As such, rewrite rules from different authors are fully interoperable.
@@
Similarly, without changing the source code, different rewrites can still occur by applying different sets of rewrite rules that can for example be offered by different Julia packages.
@@
This example also illustrates how our system supports the composition of high-level, domain-specific optimizations with optimizations of core Julia primitives. In this way, our system’s scope is much broader than recent equality saturation based optimizers such as Cranelift.
@@ 5. Compilation Time Evaluation
Julia is a just-ahead-of-time compiled language for scientific computing. By default, Julia automatically compiles a native, type-specialized version of a function just ahead of the first time it is to be executed on some combination of argument types. Julia hence positions itself somewhere between typical ahead-of-time compilers such as LLVM and gcc and just-in-time compilers such as Cranelift or the compilers used in a Java VM. Compilation speed is hence important for Julia.
@@
Alternatively, a user can request ahead-of-time compilation of functions and optionally invoke our rewriting rule-based optimization, depending on whether or not the investment in compilation time might yield a positive return on investment. As scientific software most often has relatively long execution times, compilation times are therefore not as critical as in more traditional JIT compilers. Still, we aim for interactive compilation speeds with our work to allow for rapid prototyping.
5.1. Complexity of E-Graph Construction
In general, termination for equality saturation is not guaranteed since application of rewrite rules can introduce new rewrite opportunities indefinitely. (Willsey et al., 2021; Suciu et al., 2025) In practice, this issue is most often tackled by equality saturation timeouts in combination with heuristic schedulers that ensure that all rewrite rules get a chance to be applied by implementing a rule application back-off mechanism (Cheli, 2021a; Yang et al., 2021). The non-terminating behavior of equality saturation is due to repeated application of rule application and rebuilding the e-graph by joining newly discovered equivalences. The actual construction of the e-graph, however, is not concerned with this loop and does, in fact, execute in polynomial time.
@@ 5.1. Complexity of E-Graph Construction
We reimplemented the algorithm introduced for Cranelift (Fallin, 2023) in Julia. This algorithm starts by computing the dominator tree of the input SSA code. We reuse the dominator tree construction implementation from the Julia compiler. This implementation is based on an algorithm from Georgiadis’s thesis (Georgiadis, 2005), and runs in linear time with respect to code size. IR statements are visited in a pre-order dominator tree traversal to ensure that all dependencies of the current statement have already been visited. For each statement, the e-class dependencies are looked up and an e-node is constructed. If the e-node is not part of the e-graph yet, it is inserted in a new e-class, otherwise an existing e-class is associated with the statement. At the same time that the e-graph is being constructed, the CFG skeleton is being built. This involves pushing the corresponding e-classes of effectful statements in a CFG structure which has a worst case runtime complexity of , where is the code size, but an amortized complexity of .
@@
Since no equalities are introduced yet during e-graph construction time, inserting new e-classes in the e-graph does not require running the relatively expensive rebuilding procedure. As such, e-graph construction only requires inexpensive hash map lookups and insertions for each statement, which have a worst case runtime complexity of but an expected worst case runtime of .
@@
Overall, the worst case runtime complexity of our e-graph construction is but on average and in practice, as will be discussed in Section 5.2, the worst case runtime complexity is .
@@ 5.2. Experimental Compilation Time Evaluation
A graph showing the compilation time for different parts of our rewrite optimizations. The graph consists of three vertical bar charts next to each other, representing time spent in the pre-inlining, post-inlining, and finalizations stage. Each bar chart is divided in different colors indicating the time needed for specific compilation steps. The different compilation steps are: IR to e-graph conversion, e-graph saturation, e-graph to ir conversion, type reinference, invocation queueing, ILP problem construction, ILP solver (HiGHS), optimized function insertion, inlining pass, llvm optimization and code generation, and other.
Figure 11 shows the compilation times spent in different steps of the whole optimization pipeline for optimizing the forward function from Figure 9. Important to note is that these numbers include the time spent optimizing each function in the call graph of the top-level function. For the forward method, the call graph consists of 43 methods that each pass through the full optimization pass. The chart partitions the optimization flow and its compilation time into three stages. All benchmarks were run on a machine with a 6-core Intel i7-9750H processor and 24GB RAM. The code was run on Julia v1.11.0 (commit 501a4f25c2b), with LLVM 16.0.6.
5.2.1. Pre-Inlining
In this first stage, rewrites are applied before the code body of the function is inlined. Due to the small number of rules, e-graph saturation only takes a small portion of the work in this stage. Similarly, ILP problem construction and solving runs quickly because non-inlined code consists of few statements, leading to a small ILP problem to solve. In this stage, a larger part of the run time is spent reinferring the types in the generated IR. This also includes replacing dynamic function calls with static invocations of the correct methods if enough type information is available to make this decision. The last step of this stage consists of collecting the non-inlined IR of each method that is invoked, and adding this IR to the top of a worklist. Methods invoked within a function end up at the top of the worklist and are fully optimized before the second stage of the current function can begin.
5.2.2. Post-Inlining
The second stage starts by replacing each method invocation with the optimized variant if rewrites were applied to it. Next, we have the default inlining pass that is also invoked during regular Julia compilation. In this second stage, conversion from IR to e-graph takes much longer because the inlined code consists of many more nodes. For the same reason, e-graph saturation takes longer as well. Because there are no rules to apply post-inlining in this instance, no time is spent on ILP, and type re-inference runs faster.
To assess the scalability of the conversion from IR to e-graph and back, Figure 12 shows those times for IRs of different sizes. The IR dataset was collected by running the optimizer on the factorize function from Julia’s linear algebra standard library and on the many functions that are part of this function’s call graph. Both conversions clearly scale linearly with the number of statements.
E-graph to IR conversion does not meaningfully show up in the compilation times in Figure 11 because this operation is only executed if actual changes to the e-graph were made, which happens infrequently in this instance. Similarly, re-inference times in the post-inlining stage do not take much time because this step is only applied when the e-graph has changed. This is in contrast to the pre-inlining stage, where part of the re-inference procedure needs to be executed either way to prepare the IR for inlining in the next stage.
A scatter plot showing the time needed to convert from IR to e-graph and vice versa. The x-axis shows the number of IR statements, the y-axis shows the time needed in microseconds. Datapoints for IR to e-graph conversion are shown with an orange, plus-shaped marker, and e-graph to IR is shown using an red, triangle-shaped marker. Both cases a similar linear trend upwards.
5.2.3. Finalization
During the last stage, the functions that were enqueued in the previous stage have been optimized and are inserted if rewrites were applied to them. Finally, compilation is passed on to LLVM to generate an executable, native function from the optimized IR. This step takes more than a third of the total time.
In Figure 11, Invocation Queueing, Inlining, and LLVM Compilation are steps that are not specific to our compilation flow but are also taken in the native Julia compiler. In the example discussed above, 80% of the total time is spent in these steps. The Julia compiler hides much of the compilation costs by caching the compiled code. Tighter integration of our optimization into the Julia compiler could thus lead to faster compile times as well. As briefly noted in Section 3.1, rewrite rules will be applied more often on IR before inlining because that IR contains function calls roughly as they were written by a programmer. In IR to which inlining has been applied, on the other hand, it is difficult to anticipate which function calls will end up visible, making it less productive to write rules targeting IR at this level. For this reason, a worthwhile trade-off might be to disable rewrites in the second, post-inling stage. This would eliminate the high cost of the IR to e-graph conversion that is included in the middle bar in Figure 11.
5.2.4. ILP Construction and Solving
@@
Table 2 shows the time required to construct and solve the ILP problem for several examples discussed in this paper. With the exception of the code for simplifying 2D transformations in Figure 6, e-classes do not contain many equivalences, leading to fast problem construction and solving. For the 2D transformation example, particular rewrite rules, e.g., those that combine two subsequent rotations into one rotation, give rise to an infinite number of rewrite opportunities. In these instances, saturation runs until a timeout is reached. If fast compilation times are required for pathological cases like the one shown, we can fall back to a greedy extraction algorithm. In the future, better scheduling of rules and detection of these cases could limit the amount of e-graph blow-up as well.

### v2

@@ 4.3. Multi-Line Broadcast Fusion
Julia code that uses broadcasting syntax: using the dot operator, an elementwise ReLu operation and an elementwise addition are applied to vectors A and B. In other words, the relu function is broadcasted over matrix A and vector B is added to the result using broadcasting addition.
@@
Julia’s broadcasting mechanism allows users to apply functions element-by-element on one or more arguments containing multiple elements. Syntactically, this is done by adding a dot (.) between the function and its argument list. The top of Fig. 8 shows two broadcast operations. relu.(A) applies the scalar relu function to each element of the matrix A. Next, the .+ call computes the addition between the result and a vector B. Although the two arguments to .+ have a different shape, the operation can take place because the shapes are compatible.
The Julia compiler transforms expressions containing this special dot syntax into calls to the standard library functions broadcasted and materialize. Fig. 8 shows the result for the expression in Fig. 8. A call to broadcasted builds a lazy representation of the broadcast result. Different lazy broadcast objects can be nested; it is only when materialize is called that the final result object is allocated and filled with computed values. As such, different operations within multiple broadcasted calls are fused when materialize is called. This leads to less memory being used because temporary arrays are not materialized and potentially better run-time performance because there are fewer function calls. This operator fusion is especially beneficial for code that is executed on the GPU, as kernel launches can take a significant amount of time, and because kernel fusion can significantly reduce the required number of memory accesses (Besard et al., 2019).
A limitation of Julia’s broadcasting lowering is that fusion cannot occur across broadcast expressions on different lines. This is because replacing the dot syntax with calls to broadcasted and materialize is a syntactical transformation that is performed during AST lowering, which operates statement by statement. A potential optimization is hence to eliminate superfluous calls to materialize. Rewrite rules make this easy. Take, for example, the rule This rule will remove the intermediary materialization for any call of the form f.(x) where x is a variable that was defined on a different line with a broadcast expression itself. In the example in Fig. 10, fusion will now be applied despite the two dot operators occurring in two different statements on two separate lines. This illustrates how the system we propose not only supports domain-specific optimization but also at the same time enables optimization of common code patterns involving only core Julia primitives.
An important detail is that here, once more, CFG skeleton relaxation is required, since materialize ends up calling a foreign C function which the compiler cannot deem effect-free, poisoning the remainder of the effect analysis. In practice, this means that the rule needs to be prefixed with relaxed.
@@
For other patterns, e.g., an expression where the outer broadcasted function is a binary function, separate rules need to be written. Table 1 lists how many times our analysis found different multi-line broadcasting patterns that can all be rewritten to broadcasted(~f, ~x) and broadcasted(~binop, ~x, ~y). The last three patterns are different configurations of broadcasting a binary function. Note that the matches for these three patterns are not independent: a match against the last pattern, which will result in two invocations of materialize being optimized out, implies the two other patterns can be matched and applied as well, each optimizing out only one materialize invocation. When the final code is extracted in such cases, the expression from the e-class containing the least materialize calls will be chosen.
@@ 4.4. Extending Broadcasting with Domain-Specific Batched Operations
In high-performance contexts, programmers often sacrifice code simplicity and readability in pursuit of achieving the highest possible performance. For example, instead of writing deep learning layers and non-linear activation functions using core language constructs, programmers often resort to using frameworks or libraries that contain optimized implementations of matrix multiplications, activation functions, attention computation, and others (Elshawi et al., 2021; Parvat et al., 2017). This works well for most common use cases, but falls apart when a programmer requires custom behavior that is not yet supported by the framework. In those cases, the programmer is required to rewrite parts of the program to stop using the framework and instead depend on a different framework or the core language. Instead of having to opt out of framework functionality and rewriting code, one could ideally write simple, readable code and have the frameworks themselves automatically discover opportunities to use their more performant implementations. Our system gives framework authors the option for implementing such discovery and optimization almost for free.
@@
As an example, the code in Fig. 10 shows an implementation of the forward pass of a single fully connected feedforward layer. Several existing Julia packages allow us to compute such operations more efficiently (Faingnaert et al., 2022; Innes et al., 2018; Innes, 2018). For example, the programmer could rewrite the code as shown in Fig.10 to use the optimized fused_dense_bias_activation function from the Lux.jl deep learning framework (Pal, 2023). A disadvantage of such manual rewriting is that the code becomes less readable for someone not familiar with the Lux.jl APIs. Moreover, the use of that specific framework’s optimized implementation, which is in essence a decision about a non-functional aspect of the code, is then hard-coded in the code that specifies the functionality. This requires the developers and the maintainers of this code to be knowledgeable in the code’s application domain as well as in the Lux.jl framework. Switching to other frameworks in the future will require rewriting code. Clearly, hard coding the dependence on Lux.jl and its APIs in the code has a number of downsides.
@@
With our system, instead of rewriting the original code of Fig. 10, it suffices to add a rewrite rule that maps the original code pattern onto the optimized fused_dense_bias_activation function:
| materialize(broadcasted(~, broadcasted(+, ~W::AbstractMatrix * ~x::AbstractMatrix, ~b::AbstractVector))) | ||
@@
Combined with the multi-line broadcast rewriting rule explained above, which rewrites the original code into a single line without the intermediary variable logits, this rule maps the single-line broadcast expression onto a call to the library function LuxLib.fused_dense_bias_activation, identical to what would be generated for the code in Fig. 10.
@@
This example illustrates how our approach allows for a better separation of concerns. Another advantage is that alternative rewrite rules can be provided for different Julia packages that provide different optimized implementations for similar functionality. Different rewrite rules also do not obfuscate each other. All rewrite rules can be represented in the e-graph. Only during extraction is the final optimized code materialized, optionally based on user-defined extraction costs for different functions. As such, rewrite rules from different authors are fully interoperable.
@@
Similarly, without changing the source code, different rewrites can still occur by applying different sets of rewrite rules that can, for example, be offered by different Julia packages.
@@
This example also shows how our system supports composing high-level, domain-specific optimizations with optimizations of core Julia primitives. In this way, its scope is much broader than recent EqSat optimizers like Cranelift.
4.5. Keeping up with evolving and emerging libraries for GEMMs
As an example of multiple libraries providing complementary functionality, consider General Matrix Multiplications (GEMMs), a cornerstone in many scientific domains. GEMMs are also often combined with element-wise operations or elops on their operands or result, for example with non-linear activation functions such as ReLU.
Many libraries exist to perform different GEMM variants (Van Zee and van de Geijn, 2015; Faingnaert et al., 2022; Torres et al., 2024), each offering different trade-offs and interfaces. For targeting NVIDIA GPUs, e.g., the cuBLAS library provides the highest performance for basic GEMM expressions on matrices with simple datatypes. It also offers some kernels that fuse often-used elops with the GEMM. For elops not supported by cuBLAS, the alternative GemmKernels.jl can be used (Faingnaert et al., 2022). This package leverages Julia’s multiple dispatch and GPU programming support (Besard et al., 2019) to generate performant GEMM code on the fly, including for user-defined elops and for GEMMs on less commonly used data types such as dual numbers, which are useful for automatic differentiation (Revels et al., 2016).
Over time, such libraries evolve. For example, cuBLAS has grown to include ever more kernels with fused elops. Ideally, programmers should not have to rewrite their application code to make use of newly available functionality or when switching from one library to another. Instead, they should be able to write generic code that automatically makes use of the available functionality in libraries. Rewrite rules enable this.
Take for example the expression C .= relu.( * (f_a.(A) * f_b.(B)) + * C), where A, B, and C are matrices and and are scalars. The different GEMM and linear algebra packages can each provide a custom set of rewrite rules that rewrite as much as possible of that expression into calls to optimized functions. For example, a rule added to the LinearAlgebra.jl package would allow this expression to be rewritten into a more efficient call to mul! that carries out the matrix multiplication, scaling by and , and accumulation into C at once. Then only the elops (f_a, f_b, and relu) remain to be executed in separate kernels. For particular argument types, such as 16-bit floating-point matrices in GPU memory, rewrite rules in the GemmKernels.jl package would even allow the complete expression to be fused into one single kernel. Other packages could still contain rewrite rules for other GEMM configurations or variants. Depending on which rules are loaded, the same code hence can be rewritten in different ways, exploiting the various and possibly evolving optimization opportunities that complementary libraries offer without having to rewrite the source code.
@@ 5. Compilation Time Evaluation
Julia is a just-ahead-of-time compiled language for scientific computing. By default, Julia automatically compiles a native, type-specialized version of a method just ahead of the first time it is to be executed on some combination of argument types. Julia hence positions itself somewhere between typical ahead-of-time compilers such as LLVM and gcc and just-in-time compilers such as Cranelift or those used in a Java VM. Compilation speed is hence important for Julia.
@@
A user can also request ahead-of-time compilation of methods and optionally invoke our rewriting rule-based optimization, based on whether or not the investment in compilation time might yield a positive return on investment. As scientific software often has relatively long execution times, compilation times are hence not as critical as in traditional JIT compilers. Still, we aim for interactive compilation speeds with our work to allow for rapid prototyping.
5.1. Complexity Analysis of Equality Saturation
E-Graph Construction
@@ 5.1. Complexity Analysis of Equality Saturation
We reimplemented the algorithm that converts IR into an e-graph introduced for Cranelift (Fallin, 2023) in Julia. This algorithm starts by computing the dominator tree of the input SSA code. We reuse the dominator tree construction implementation from the Julia compiler. This implementation is based on an algorithm from Georgiadis’s thesis (Georgiadis, 2005), and runs in linear time with respect to code size. IR statements are visited in a pre-order dominator tree traversal to ensure that all dependencies of the current statement have already been visited. For each statement, the e-class dependencies are looked up and an e-node is constructed. If the e-node is not part of the e-graph yet, it is inserted in a new e-class, otherwise an existing e-class is associated with the statement. At the same time that the e-graph is being constructed, the CFG skeleton is being built. This involves pushing the corresponding e-classes of effectful statements in a CFG structure which has a worst-case execution time complexity of , where is the code size, but an amortized complexity of .
@@
Since no equalities are introduced yet during conversion from IR to e-graph inserting new e-classes in the e-graph does not require running the relatively expensive e-graph rebuilding procedure. As such, e-graph construction only requires inexpensive hash map lookups and insertions for each statement, which have a worst-case execution time complexity of but an expected worst-case execution time of .
@@
Overall, the worst-case execution time complexity of our e-graph construction is but on average and in practice, as will be discussed in Section 5.2, the worst-case execution time complexity is .
E-Graph Saturation
In general, termination for EqSat is not guaranteed because the application of rewrite rules can introduce new rewrite opportunities indefinitely (Willsey et al., 2021; Suciu et al., 2025). In practice, this issue is most often addressed by EqSat timeouts in combination with heuristic schedulers that ensure that all rewrite rules get a chance to be applied by means of a rule application back-off mechanism (Cheli, 2021a; Yang et al., 2021). The non-terminating behavior of EqSat is due to repeated rule application and rebuilding the e-graph by joining newly discovered equivalences.
Extraction
Constructing the ILP problem entails a few different steps. First, simple cycles are determined using Johnson’s algorithm (Johnson, 1975), and the dominator tree is constructed. Since the number of cycles in a graph can grow exponentially in the number of nodes (Johnson, 1975), in the worst case, the time required to construct the problem is exponential as well. In practice, however, e-graphs typically contain only a few short simple cycles. Next, the minimization objective itself can be constructed in , with the number of e-nodes, and the number of statements in the CFG skeleton. Assuming a low number of class cycles, the problem construction time is asymptotically dominated by the construction of the class constraint (Equation 3). In the worst case, this takes time, with the maximum number of children e-classes of an e-node, the maximum number of dominating statements, and the maximum number of e-nodes in an e-class.
Solving the ILP problem is an NP-hard problem, but as will be shown in Section 5, solving times are of the same order of magnitude as construction.
@@ 5.2. Experimental Compilation Time Evaluation
We conducted our measurements on a 64-core AMD Ryzen Threadripper PRO 7985WX and 502GiB RAM, with Julia Version 1.11.5 (LLVM 16.0.6). Fig. 11 shows the time taken for the different steps of our optimization scheme for Fig. 10’s forward function (invoked with training=false) and the rewrites to fuse the operations as discussed in Section 4.4.
Conversion
E-graph conversion, both from and to Julia IR, takes relatively little time for this example. To evaluate the run time of conversions for other code, Fig. 12 shows the time needed to convert between IR and E-Graph for many different Julia methods. These methods were selected arbitrarily by collecting all the methods in the call graphs encountered during the compilation of some methods from Julia’s core and linear algebra standard library. While our optimization targets IR without inlined function calls, we also include timing results for IR with inlined calls to capture performance on longer IR. Fig. 13 shows that the vast majority of the evaluated methods consists of only a handful of IR statements. This means that we typically do not expect large e-graphs during optimization.
Code Generation
Not shown in the compilation time overview is the time that LLVM spent to optimize and generate the final executable code from Julia IR. The reason for the omission in the figure is that this step also occurs for regular Julia code compilation. For the example from Fig. 11, LLVM code generation takes around 75ms, or about 95% of total compilation time. For regular Julia code, much of the compilation cost imposed by LLVM is hidden by caching of compiled code. Tighter integration of our optimization into the Julia compiler could similarly help hide this cost as well.
ILP Construction and Solving
@@
Fig. 11 shows that most of the time is occupied by constructing and solving the ILP extraction problem presented in Section 3.5. Apart from the inherent computational complexity of solving the ILP problem, the conversion takes a large amount of time for converting the e-graph into a collection of variables and constraints in an external ILP solver library. Table 2 shows the time required to construct and solve the ILP problem for several use cases discussed in this paper. With the exception of the code for simplifying 2D transformations in Fig. 6, e-classes do not contain many equivalences, leading to fast problem construction and solving. For the 2D transformation example, particular rewrite rules, e.g., those that combine two subsequent rotations into one rotation, give rise to an infinite number of rewrite opportunities. In these instances, saturation runs until a timeout is reached. If fast compilation times are required for pathological cases like the one shown, we can fall back to a greedy extraction algorithm. In the future, better scheduling of rules and detection of these cases could limit the amount of e-graph blow-up as well.
| Construction [ms] | Solve [ms] | Variables | Constraints | E-Classes | E-Nodes | |
| Fig. 4 | 0.46 | 1.042 | 126 | 364 | 26 | 28 |
| Fig. 6 | 24.16 | 208.66 | 918 | 5325 | 337 | 1,002 |
| Fig. 10 | 0.37 | 5.02 | 48 | 168 | 21 | 23 |
6. Performance Evaluation
| Use Case | Description | Speed-up |
| Fig. 6 | 2D transformations | |
| Fig. 4 | Linear algebra value reuse | up to |
| Fig. 10 | Fused neural network | up to |
| Sec. 4.5 | GEMM rewrites | up to |
Our work does not at all aim to squeeze more performance out of the Julia programming language and its compiler and run-time implementation. Indeed, any optimization that can be obtained with a rewrite rule can also be obtained by manually rewriting source code, such as by duplicating source code and then special-casing it. Still, as the goal of the rewrite rules is to enable optimizations, this section reports the performance gains that can be achieved for some of the previously presented use cases.
Our results are summarized in Table 3. The use case of Fig. 6 where domain-specific code representing a 2D transformation is optimized into an idempotent function shows a large speed-up, as expected by all computations being optimized out of the function. The other examples paint a more interesting picture.
For the use case of Fig. 4, where the computation of a matrix determinant is simplified by reusing the already computed eigenvalues, Fig. 14 shows a moderate speed-up across the board. For problem sizes around , this speed-up is more pronounced. For reference, yellow dots show the speed-up when simply omitting the determinant computation completely. In other words, they show what speed-up can be achieved by reducing the execution time of the determinant computation to zero. The fact that those roughly match with the optimized code shows that in the rewritten code, the determinant value basically comes for free, whereas it costs up to 10% in the original code.
Similarly, for the use case of Fig. 10 where the forward pass of a neural network is replaced with a fused implementation, the plot in Fig. 16 shows a speed-up for small problem sizes but ends up about as fast as the original code for larger problem sizes, as the time spent in the matrix multiplication tends to dominate. 222For problems involving GPU computation (Fig 16, Fig. 16), we measured execution time on a machine with a 16-core Intel Xeon E5-2637 v2, NVIDIA GeForce RTX 2080 Ti (12GB), and 64GB RAM, with Julia Version 1.11.5 (LLVM 16.0.6). For the other problems, we use the same setup as in Section 5.
Finally, we look at the example of rewriting the expression C .= relu.( * (f_a.(A) * f_b.(B)) + * C), discussed in Section 4.5. Fig. 16 shows that the basic linear algebra rewrite, which fuses everything except the elops (f_a, f_b, relu) into one GPU kernel. This provides a moderate speed-up across the board over the baseline implementation in which each operation in the whole expression corresponds to a separate GPU kernel. The GemmKernels rewrite, which fuses all operations into a single kernel, is worthwhile for smaller problem sizes, and shows a small slowdown for larger problems, where the core GEMM computation dominates.
This performance evaluation of four uses cases demonstrates that our system can indeed be used to achieve significant speed-ups by exploiting domain-specific optimization opportunities, in a programmer-friendly manner.